OpenStudy (anonymous):

What polynomial has roots of –1, –2, and 3

OpenStudy (saifoo.khan):

x = -1 x = -2 x = 3

OpenStudy (saifoo.khan):

(x+1) = 0 (x +2) = 0 (x -3) = 0 Finally, (x+1)(x+2)(x-3) = 0

OpenStudy (saifoo.khan):

Multiply! [:

OpenStudy (saifoo.khan):

solved?

OpenStudy (anonymous):

x3 + 6x2 + 11x + 6 x3 – x2 – 5x + 2 x3 + 2x2 – 7x + 2 x3 – 7x – 6 these are the answers

OpenStudy (saifoo.khan):

Try to solve it too!!!!!!!!!

OpenStudy (saifoo.khan):

LOL.

OpenStudy (saifoo.khan):

I'm not a calculator.

OpenStudy (saifoo.khan):

Jk.

OpenStudy (anonymous):

ok

OpenStudy (saifoo.khan):

\[\Huge x^3-7 x-6 = 0 \]

OpenStudy (anonymous):

but that not one of the answers

OpenStudy (saifoo.khan):

LOL, what time is it there??

OpenStudy (saifoo.khan):

Grab a soda!! O_O it will open your eyes! ^.^ JK.

OpenStudy (saifoo.khan):

it's D part.

OpenStudy (anonymous):

oh my gosh it is only 630pm here where r u

OpenStudy (anonymous):

ok so r u ready for more or u r done with m lol?

OpenStudy (saifoo.khan):

lol, im from Pakistan. It;s 03:24am here.

OpenStudy (anonymous):

omg what r u doing up so early,? but i am so glad that u r up

OpenStudy (saifoo.khan):

the thing is, i have to run to bed. LOL. so if u have more u may post them asap, so u can wait untill tmmrw!!

OpenStudy (saifoo.khan):

LOL, i didn't slept yet. Im up from whole day!! xD